OpenStudy (anonymous):

If h(x) = x − 1 and j(x) = −4x, find h[j(5)].

OpenStudy (anonymous):

h(j(x))=-4x-1 and h(j(5))=-4(5)-1=-21

OpenStudy (anonymous):

Thank you . : )